OpenStudy (anonymous):

Express the complex number in trigonometric form. -3i 3(cos 180° + i sin 180°) 3(cos 270° + i sin 270°) 3(cos 90° + i sin 90°) 3(cos 0° + i sin 0°)

OpenStudy (anonymous):

@arose114 I believe it's C.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

@GLKK you're not suppose to give straight answers people ask questions because they don't understand something please don't give answers without explaining or walk them through so they can engage in solving

jimthompson5910 (jim_thompson5910):

the answer is wrong anyway

OpenStudy (jdoe0001):

well... notice the coordinates \(\large { -3i\implies 0-3i\implies (x,y) \\ \quad \\ \begin{array}{llll} 0&-3\\ x&y \end{array} }\) graph that... what angle does it give you?

OpenStudy (anonymous):

270

OpenStudy (anonymous):

so B?

OpenStudy (anonymous):

http://www.ditutor.com/complex_numbers/trigonometric_form.html heres a website that might help understanding

OpenStudy (jdoe0001):

yeap 270 and the "r" is well... |-3| so 

OpenStudy (anonymous):

Ahh okayy!!! Thank you so much!

OpenStudy (jdoe0001):

r is really just the hypotenuse of x, y so \(\bf r=\sqrt{x^2+y^2}\implies r\sqrt{0+3^2}\to 3\)

OpenStudy (jdoe0001):

yw